lisa has a piece of polka- dot ribbon that is as long as 5 pieces of 1 1/4. how long is lisas polka-dot ribbon .
Answer by stanbon(75887)

lisa has a piece of polka- dot ribbon that is as long as 5 pieces of 1 1/4. how long is lisas polka-dot ribbon
-------
Ans: length = 5*(1 1/4) = 5(5/4) = 25/4 = 6 1/4
-------------------
Cheers,
Stan H.
===================